THE DUAL OF THE INVARIANT QUINTIC

EBERHARD FREITAG AND BRUCE HUNT

Abstract. We have identified explicitly the dual of the hypersurface I5 defined as the unique quintic hypersurface in P^{5}ℂ invariant under the natural action of the Weyl group W (E_{6}) of the lattice E_{6}.